# Magic tricycles: efficient magic state generation with finite block-length quantum LDPC codes

### v1

@@ 2.2 Transversal $CCZ$ circuits
    $f_{CCZ}$ must vanish when any argument corresponds to a binary string $b$ defining an $X$-type stabilizer – i.e $X^{b}=X^{b_{1}}\otimes X^{b_{2}}\cdots\otimes X^{b_{3n_{G}}}$ is an X-type stabilizer. This condition ensures $f_{CCZ}$ descends to a well-defined function on the space of logical codewords.
@@
We expand on both conditions and their homological interpretation in Appendix˜B. See Proposition 2 for how $f_{CCZ}$ is defined for tricycle codes.
Breuckmann et al. [16] introduced a general formalism for constructing appropriate trilinear functions $f_{CCZ}$ that satisfy conditions $1$ and $2$. Using this framework, we show that all tricycle codes with check-weights $(w_{X},w_{Z})=(12,8)$ admit depth-18 transversal $CCZ$ circuits. Shallower circuits – down to depth 4 – can be constructed by imposing additional structure on the code, such as constraints on the matrices $\mathbf{A},\mathbf{B},\mathbf{C}$ used to define the parity checks in Eq.˜1. Details of these constructions appear in Appendix˜B, and representative examples are listed in Table˜2. While shorter-depth codes tend to have lower distance or rate, we observe that increasing block size recovers performance, suggesting no intrinsic limit on code parameters from enforcing such circuit-depth reducing conditions.
We distinguish two classes of transversal circuits:
  • Single-sector circuits, where each $CCZ$ gate acts on one sector per block – e.g., sector $I$ of block 1, $II$ of block 2, and $III$ of block 3 (illustrated in Fig.˜1b).
  • Triple-sector circuits, where all sectors of all blocks participate in a layered $CCZ$ pattern. These are organized into parallelizable layers for constant depth (see Fig.˜1c).
@@
Using the methods described in Appendix˜C, we are able to find at-least $K_{CCZ}\geq 5$ for the $N=192,375,405,648$ qubit codes and $K_{CCZ}\geq 4$ for the $N=72,144$ codes in footnote˜2 using the depth $18$ CCZCCZ circuit – see Appendix˜B. We attempted to find larger values of $K_{CCZ}$ for these codes, but the solver we used did not converge within a reasonable amount of time. We therefore emphasize that the $K_{CCZ}$ values found here are lower bounds returned by exact methods that quickly become infeasible for larger $K_{CCZ}$ – this is primarily due to the current lack of well-performing efficient heuristics for the binary tensor-subrank problem, and we anticipate that progress towards such heuristics will lead to a higher yield of individually distillable $\overline{CCZ}$ type magic states for the codes in footnote˜2.
On the other hand, for some of the codes with short-depth $CCZ$ circuits from LABEL:table::short_code_polys, we were able to determine the exact values of $K_{CCZ}$. For the $N=135$ codes with depth $4$ and depth 88 CCZCCZ circuits from LABEL:table::short_code_polys, we found $K_{CCZ}=1$ in both cases. For the $N=405$ depth $4$ and depth $8$ codes, we similarly found $K_{CCZ}=3$ in both cases. A more extensive search over tricycle codes with shorter $CCZ$ circuit depths may reveal codes with larger $K_{CCZ}$ – our search over codes with short-depth circuits is far from exhaustive, and we did not select for short-depth codes with large $K_{CCZ}$ in the numerical search, focusing instead on code parameters. However, it remains a possibility that imposing the additional constraints needed to reduce the depth of $CCZ$ circuits (see Appendix˜B) generically leads to lower attainable $K_{CCZ}$ values – further work is needed to ascertain whether this is the case. For completeness, we note that for the depth $12$ and $16$ CCZCCZ circuit codes in Table˜2, our methods once again returned lower bounds of $K_{CCZ}\geq 4$.
$\mathbf{[[N,K,(D_{X},D_{Z})]]}$ CCZ depth Circuit type
$135,3,(\leq 22,8)]]$ 4 $S$
$405,9,(\leq 27,10)]]$ 4 $S$
$135,3,(\leq 27,9)]]$ 8 $S$
$405,9,(\leq 29,11)]]$ 8 $S$
$405,9,(\leq 37,14)]]$ 12 $S$
$405,9,(\leq 27,11)]]$ 12 TS
$375,15,(\leq 15,10)]]$ 16 TS
Table 2: Examples of codes with code-space preserving CCZ circuits with depths less than 18. The column labeled ‘Circuit type’ indicates whether the associated circuit is a single-sector ($S$) or triple-sector (TS) circuit. See Appendix˜B and LABEL:table::short_code_polys for details of how these codes are constructed.
@@ 2.3 Single-shot distillation
However, initial $Z$-type stabilizer measurements produce non-deterministic $\pm 1$ outcomes, which must be reliably fixed to +1+1 on hardware before applying the $CCZ$ gate [5]. In contrast to codes like the surface code, which lack single-shot state preparation, tricycle codes embed a sufficient set of metachecks—additional redundancies that allow a decoder to identify and correct syndrome measurement errors. Metachecks essentially form a separate code that protects the stabilizer syndromes themselves. For the $Z$ basis, these take the form
@@
where the $\mathbf{A},\mathbf{B},\mathbf{C}$ matrices are the same as in Eq.˜1. The tolerance to measurement qubit errors is determined by the single-shot distance $D_{Z}^{SS}$ [20], which is defined as the minimum weight of a faulty syndrome that passes all meta-checks but is not a $Z$ syndrome. A large $D_{Z}^{SS}$ indicates considerable redundancy, facilitating the identification of sparse syndrome errors. For all codes examined (see footnote˜2 and Table˜2), we find $D^{SS}_{Z}=D_{Z}$, or at least share their upper bounds as obtained by Monte Carlo methods (see Appendix˜D).
@@
In conjunction with soundness properties [20] and under select error models, metachecks have been shown to allow for a decoder to fault-tolerantly infer the $Z$ stabilizer eigenvalues after only a constant number of error correction rounds. It is worth noting that this task of preparing the $Z$ stabilizers to +1+1 from a completely random initialization is a more challenging task than just single shot error correction [43, 7], where a decoder may assume that the $Z$ stabilizers were fault-tolerantly fixed to +1+1 prior to the observed noise. For example, many hypergraph product codes enable single-shot error correction but not single-shot state preparation [43].
@@
We supplement our theoretical understanding of the tricycle codes’ single-shot state preparability with a direct simulation of single-shot state preparation for three codes with good finite-size encoding rates and distances. Our results, shown in Fig.˜3 demonstrate exponential suppression of the logical error with increased distance, consistent with fault-tolerant state preparation.
@@
Figure 3: Phenomenological noise simulation of single-shot state preparation in the $Z$ basis. Our method resembles Ref. [43] and assumes that the initial $Z$ syndrome is trivial. We apply a round of measurement errors with probability $p$, followed by two more rounds of syndrome measurement. Pauli $X$ errors occur in each on the data qubits with probability $p$, in addition to measurement errors. The three rounds of syndrome data are jointly decoded by a BP+OSD decoder [72], corresponding to a three-round windowed decoding protocol (Sec.˜2.4). After the inferred correction operator is applied, we perform a noisy transversal $Z$ basis measurement of the data qubits, decode the reconstructed syndrome, and apply the corresponding correction. A logical failure is said to occur if the residual $X$ operator is a logical operator of the tricycle code. Logical error rates are determined via Monte Carlo simulations; error bars indicate standard errors, computed as $\sqrt{p_{L}(1-p_{L})/M}$, where $M$ is the number of samples.
@@
Going back to the magic state factory, following initialization, the single-shot error correction capabilities of tricycle codes are paramount to prevent errors from spreading. Each $CCZ$ gate will propagate a pre-existing $X$ error into $CZ$ errors on the other two qubits, which are collapsed into nondeterministic patterns of $CZ$ errors after syndrome measurement. To prevent such errors from building up, the $Z$ stabilizers must continuously be repaired to +1+1. In codes without single-shot properties, this usually requires ‘just-in-time decoding’ [8], in which error clusters are allowed to expand modestly before being corrected, and accompanies a decrease in performance and threshold of the code [75, 76]. The single shot $Z$ basis of the tricycle code, in contrast, is able to immediately correct $X$ errors (up to a small residual error), limiting repeated propagations of $CZ$ errors. Although the $X$ basis is not single-shot, importantly, the $Z$ errors they detect commute with the $CCZ$ gates, so they do not need be corrected in real time. This is opportune given that the $CCZ$ circuit may disrupt the $X$ stabilizer group until the circuit is completed. As the $CCZ$ circuit is of constant depth, it suffices to measure the $X$ stabilizers only after completion; however, incorporating intermediate correction strategies that leverage stabilizer masking [60] could be advantageous.
@@
We emphasize that these techniques for single-shot distillation are highly unoptimized, and significant improvements in performance are plausible. For instance, post-selecting factory outputs on soft information – the confidence level of the decoder in the produced states’ correctness – is known to considerably decrease the probability of logical failure with only a slight resource overhead [34, 10, 4, 5]. We leave the optimization of this single-shot distillation scheme to future work.
@@ 2.4 Circuit-noise simulations
We now evaluate the performance of our codes under realistic circuit-level noise models. Specifically, we demonstrate robust single-shot error correction in the $Z$ basis and assess fault-tolerant behavior in the $X$ basis using repeated rounds of syndrome extraction [13].
@@
To benchmark performance, we simulate errors affecting the entangling gates within the syndrome extraction circuits of tricycle codes, adopting a standard two-qubit depolarizing error model. For a given two-qubit gate error rate $p_{2q}$, each entangling gate is subject, with equal probability, to one of the fifteen nontrivial two-qubit Pauli errors from $IX,IY,\dots,ZZ$}, and experiences no error with probability $1-p_{2q}$.
We perform two types of simulations: (1) single-shot simulations in the $Z$ basis, and (2) conventional $d$-round simulations in the $X$ basis. For the single-shot scenario, we utilize a windowed decoding protocol, wherein each window comprises a fixed number of syndrome extraction rounds followed by decoding and correction. This window is then repeated multiple times, providing a concrete probe of the code’s sustainable logical error suppression [18]. Specifically, our simulations employ windows of three rounds, each repeated fourteen times for a total of forty-two rounds—a protocol previously demonstrated to saturate sustainable performance in high-rate codes [95]. With fixed window size, only codes possessing genuine single-shot capability exhibit logical error suppression that scales with code distance; in contrast, codes lacking such features see their performance limited by the window size. Although a window size of one is theoretically sufficient, empirical studies indicate that size three yields enhanced robustness to measurement errors and generally improved logical error rates [95, 12]. For conventional $d$-round simulations in the $X$ basis, the syndrome extraction is repeated for $d$ cycles, and decoding is performed using the entire record of accumulated syndrome information.
Decoding is achieved by recasting the syndrome extraction circuits as spacetime codes: checks correspond to parities of stabilizer measurements across consecutive time frames, and qubits are associated with distinct fault locations in the circuit [95, 12]. We employ belief propagation with ordered statistics decoding (BP+OSD) [72] to process the observed syndrome data. In single-shot simulations, belief propagation is applied to each window, except for the final readout which uses BP+OSD; for $d$-round simulations, BP+OSD is used for decoding the full syndrome record. The $d$-round simulations are computationally intensive, particularly for larger codes such as the $375,15,\leq 15]]$ and $648,18,\leq 18]]$ instances. Additional details, including decoder parameters and simulation protocols, are provided in Appendix˜E.
Our simulation results are summarized in Fig.˜4. Logical error rate is defined as the probability that any logical qubit within the code block fails under the simulated noise model. We observe a gate-noise threshold greater than $0.4\%$ for each of the $X$ and $Z$ basis results (see the discussion after Theorem˜A.1 in for further details). By fitting the measured logical error rates to an exponential decay model, $p_{L}=\alpha(p/p_{\text{th}})^{\beta n^{\gamma}}$, we estimate the achievable error rates for both $X$ and $Z$ bases. Table˜3 presents logical error rates for two key physical error benchmarks: $p_{2q}=10^{-3}$, representative of near-term hardware, and $10^{-4}$, representative of projected advances, together with relevant code parameters.
@@
Figure 4: Circuit-level noise simulation results for tricycle codes. (a) Logical error rate (as measured by block failure probability) versus two-qubit physical gate error rate ($p_{2q}$) for single-shot error correction in the $Z$ basis. Single-shot performance is evaluated using a windowed decoding protocol: three rounds of syndrome extraction followed by decoding and correction, with the window repeated 14 times (for 42 total rounds) to probe sustainable suppression of logical errors. (b) Logical error rate versus $p_{2q}$ for fault-tolerant, $d$-round error correction in the $X$ basis. In both panels, errors are sampled under a standard two-qubit depolarizing circuit-level noise model, and results are shown for various tricycle code. Logical error rates are determined via Monte Carlo simulations using a BP+OSD decoder, with each data point corresponding to $M$ samples; error bars indicate standard errors, computed as $\sqrt{p_{L}(1-p_{L})/M}$.
Table 3: Example tricycle codes and their simulated logical error rates under circuit-level noise. Logical error rates $p_{L}^{(X)}$ and $p_{L}^{(Z)}$ are shown for both $p_{2q}=10^{-3}$ and $p_{2q}=10^{-4}$, for different [[N, K, D]] code parameters.

### v2

@@ 2.2 Transversal $CCZ$ circuits
    $f_{CCZ}$ must vanish when any one or more of the arguments correspond to a binary string $b$ defining an $X$-type stabilizer – i.e $X^{b}=X^{b_{1}}\otimes X^{b_{2}}\cdots\otimes X^{b_{3n_{G}}}$ is an X-type stabilizer – and the other arguments similarly correspond to non-trivial X-type logical operators. This condition ensures $f_{CCZ}$ descends to a well-defined function on the space of logical codewords.
@@
We expand on both conditions and their homological interpretation in Appendix˜D. See Proposition 5 for how $f_{CCZ}$ is defined for tricycle codes.
In recent work, Breuckmann et al. [18] introduced a formalism for constructing appropriate trilinear functions $f_{CCZ}$ that satisfy conditions $1$ and $2$ based on algebraic operations on homological product codes known as cup-products. We develop a modification of this formalism and derive conditions that tricycle codes must satisfy to host such transversal $CCZ$ circuits – see Appendix˜D for details on the conditions. For reasons elaborated in Appendix˜D, we refer to this framework and the associated conditions on the codes as the symmetric triple cup-product formalism for transversal $CCZ$ gates, which applies to both three-dimensional homological product and balanced product [19] quantum LDPC codes – the latter of which tricycle codes are an example of. We then show that there are tricycle codes with favorable rates and distances which satisfy the conditions of the symmetric cup-product framework. We note that tricycle codes constructed using the original cup-product conditions of [18] have recently been studied in [51, 60] – the resulting 2222-2-2 codes always have $K=3$ while 4224-2-2 and other higher weight codes have $D=2$. We find that the new symmetric cup-product conditions we derive in this work are necessary to circumvent these limitations, and are more generally applicable to other classes of higher-dimensional balanced product quantum codes.
In the following discussion, as a slight abuse of terminology, we refer to the ‘depth’ of the $CCZ$ circuit and the maximum degree of any physical qubit – i.e the number of $CCZ$ gates a qubit is involved in – interchangeably. Indeed the maximum degree is the relevant quantity that controls the spread of errors between physical qubits of the code-blocks. The minimal depth for scheduling the gates in the circuit is upper bounded by a small constant factor of the degree (see Appendix˜D) [75].
The 4224-2-2 type tricycle codes which satisfy the conditions of the symmetric triple cup-product generically host depth 88 CCZCCZ circuits. Similarly, the 4424-4-2 codes have depth $16$ or depth $32$ circuits, and the 4444-4-4 codes have depth $12$, $64$, or $128$ depth circuits. We explain the choices made in the code-construction that lead to these circuit depths in Appendix˜D.
In addition to the symmetric triple cup-product framework, we introduce a more general formalism for constructing valid $f_{CCZ}$ functions on balanced-product quantum codes [19] that is amenable to an extensive numerical search method for codes with block-lenghts of a few hundred qubits – see Appendix˜E for details. Using this numerical method, we are able to find additional 4444-4-4 tricycle codes with better rates and distances along with shorter-depth physical $CCZ$ circuits than those that result from the cup-product formalism – we give two examples of such 4444-4-4 codes in footnote˜2, while we leave a more thorough exploration of this method for future work.
@@
Using the methods described in Appendix˜F, we are able to find at-least $K_{CCZ}\geq 2$ for all the codes in footnote˜2 using the associated $CCZ$ circuits. We attempted to find larger values of $K_{CCZ}$ for these codes, but the solver we used did not converge within a reasonable amount of time. We therefore emphasize that the $K_{CCZ}$ values found here are lower bounds returned by exact methods that quickly become infeasible for larger $K_{CCZ}$ – this is primarily due to the current lack of well-performing efficient heuristics for the binary tensor-subrank problem, and we anticipate that progress towards such heuristics will lead to a higher yield of individually distillable $\overline{CCZ}$ type magic states for the codes in footnote˜2.
@@ 2.3 Single-shot distillation
However, initial $Z$-type stabilizer measurements produce non-deterministic $\pm 1$ outcomes, which must be reliably fixed to +1+1 on hardware before applying the $CCZ$ gate [7]. In contrast to codes like the surface code, which lack single-shot state preparation, the $Z$ parity check matrices for tricycle codes contain a large number of redundant checks. These subsequently form a robust set of metachecks, which resolve a classical code on the syndrome bits, allowing a decoder to identify and correct syndrome measurement errors. In the $Z$ basis, these take the form
@@
where the $\mathbf{A},\mathbf{B},\mathbf{C}$ matrices are the same as in Eq.˜1. The tolerance to measurement qubit errors is determined by the single-shot distance $D_{Z}^{SS}$ [22], which is defined as the minimum weight of a faulty syndrome that passes all metachecks but is not a $Z$ syndrome. A large $D_{Z}^{SS}$ indicates considerable redundancy, facilitating the identification of sparse syndrome errors. For all codes examined (see footnote˜2), we found $D^{SS}_{Z}=D_{Z}$, or at least their upper bounds when exact distances were intractable to compute (see Appendix˜C). Concurrently, $D^{SS}_{Z}=D_{Z}$ was proven for this $H_{meta}$ analytically in [51]. It is also always possible to extend $H_{meta}$ into a full-rank matrix [22] by forcing the decoder to repair the noisy syndrome into one that can be produced by a pattern of data qubit errors, resulting in $D^{SS}=\infty$.
@@
Metachecks combine with soundness properties [22] to enable fault-tolerant inference of the $Z$ stabilizer eigenvalues after only a constant number of error correction rounds under select error models. It is worth noting that this task of preparing the $Z$ stabilizers to +1+1 from a completely random initialization is a more challenging task than just single shot error correction [47, 9], where a decoder may assume that the $Z$ stabilizers were fault-tolerantly fixed to +1+1 prior to the observed noise. For example, many hypergraph product codes enable single-shot error correction but not single-shot state preparation [47].
@@
In Appendix˜B, we prove soundness properties for a certain sub-class of tricycle codes in the $Z$ basis and review the ensuing guarantee on single-shot $Z$ basis decodability, which applies to state preparation [109]. Moreover, we perform direct simulations of single-shot state preparation for three codes with good finite-size encoding rates and distances, showing that general tricycle codes appear to host single-shot state-preparation properties. Our results, shown in Fig.˜3 demonstrate exponential suppression of the logical error with increased distance, consistent with fault-tolerant state preparation.
@@
Figure 3: Phenomenological noise simulation of single-shot state preparation in the $Z$ basis for 4224-2-2 tricycle codes. Our method follows Ref. [47] and assumes that the initial $Z$ syndrome is trivial. For each code, we simulate one round of syndrome measurement in which measurement errors occur with probability $p$, though we expect performance to improve with a larger decoding window (see Sec.˜2.4). A most likely error (MLE) decoder applies a minimum weight correction to both the data and measurement qubits. Then, we simulate a noisy transversal $Z$ basis measurement of the data qubits, decode the reconstructed syndrome with the MLE decoder, and apply the corresponding correction. A logical failure is said to occur if the residual $X$ operator is a logical operator of the tricycle code, and the logical error rate is normalized per logical qubit. The observed phenomenological threshold is $\gtrsim 13\%$. Logical error rates are determined via Monte Carlo simulations; error bars indicate standard errors, computed as $\sqrt{p_{L}(1-p_{L})/M}$, where $M$ is the number of samples.
@@
Going back to the magic state factory, following initialization, the single-shot error correction capabilities of tricycle codes are necessary to prevent errors from spreading. Each $CCZ$ gate will propagate a pre-existing $X$ error into $CZ$ errors on the other two qubits, which are collapsed into nondeterministic patterns of $Z$ errors after syndrome measurement. To prevent such errors from building up, the $Z$ stabilizers must continuously be repaired to +1+1. In codes without single-shot properties, this usually requires ‘just-in-time decoding’ [10], in which error clusters are allowed to expand modestly before being corrected, and accompanies a decrease in performance and threshold of the code [86, 87]. The single shot $Z$ basis of the tricycle code, in contrast, is able to immediately correct $X$ errors (up to a small residual error), limiting repeated propagations of $CZ$ errors. Although the $X$ basis is not single-shot, importantly, the $Z$ errors they detect commute with the $CCZ$ gates, so they do not need be corrected in real time. This is opportune given that the $CCZ$ circuit may disrupt the $X$ stabilizer group until the circuit is completed. As the $CCZ$ circuit is of constant depth, it suffices to measure the $X$ stabilizers only after completion; however, incorporating intermediate correction strategies that leverage stabilizer masking [69] could be advantageous.
@@ 2.4 Circuit-noise simulations
We now evaluate the performance of our codes under realistic circuit-level noise models. Our focus is on the 4224-2-2 family of tricycle codes, characterized by modest check weights ($w_{x}=8$, $w_{z}=6$) and featuring the shortest-depth $CCZ$ circuits from the symmetric triple cup-product construction (Appendix˜D). Circuit-level noise simulations of certain 4444-4-4 codes are presented in Appendix˜G for comparison.
@@
To benchmark performance, we simulate circuit-level noise affecting the entangling gates in the syndrome extraction circuits. We adopt a standard two-qubit depolarizing noise model: for a given two-qubit gate error rate $p_{2q}$, each entangling operation is followed, with equal probability, by one of the fifteen nontrivial two-qubit Pauli errors from IX,IY,,ZZ{IX,IY,...,ZZ}, and otherwise experiences no error with probability $1-p_{2q}$.
Tricycle codes naturally exhibit higher distance in the $X$ basis compared to the $Z$ basis. Here, we prioritize simulating the more challenging $X$-basis. $Z$-basis results, demonstrating robust single-shot behavior, are included in Appendix G. In the $X$-basis, we use the conventional $d$-round protocol: the syndrome extraction sequence is repeated for $d$ cycles, and decoding is performed using the full record of accumulated syndrome information from all rounds. We utilize the depth-optimal syndrome extraction circuit structure outlined in Sec.˜2.5.
For decoding, we map the syndrome extraction process to a spacetime code, where the checks correspond to parities of stabilizer measurement outcomes across consecutive time steps, and each code qubit is associated with a distinct fault location in the circuit [107, 14] (see Appendix˜G).
We perform the decoding for the $d$-rounds simulation of the 4224-2-2 codes using a most-likely error (MLE) decoder. The resulting logical error rates, as shown in Fig.˜4(b), are defined as the probability of a logical qubit failure, normalized both by the total number of error correction rounds and by the number of logical qubits in the block. We observe a gate-noise threshold of approximately $0.5\%$, and achieve low logical error rates at physically relevant gate errors. For example, we achieve a logical error rate of $\sim 4\times 10^{-5}$ at a physical error rate of $0.1\%$ for the $108,6,(d_{z}=6,d_{x}=12)]]$ code.
For further performance gains, we consider post-selection, which is standard in the magic-state distillation factory setting [16]. Specifically, we analyze the $[[48,6,(4,8)]]$ code decoded using BP+LSD [46] in conjunction with the clustering post-selection metric introduced in Ref. [57]. After $d_{z}=4$ rounds of error correction, the logical error rates as a function of abort rate using the $\alpha=2$ clusters llrs” post-selection metric of Ref. [57] are depicted in Fig.˜4(a). Additionally, we assess full error detection whereby we discard any samples that have any stabilizers flipped. We observe that for the $48$-qubit code, no stabilizer flip in roughly $32\%$ of samples, corresponding to an acceptance fraction of $32\%$ and yielding a very low logical error rate near $6\times 10^{-10}$. For the $84$-qubit code the acceptance fraction for full error detection over $d$ rounds becomes $\sim 8.5\%$ with a logical error rate $7\times 10^{-14}$.
In summary, the 4224-2-2 tricycle codes demonstrate strong performance in the $d$-round setting under circuit-level noise. It is worth emphasizing that in practical settings, $d$ rounds may not always be necessary — e.g., when transversal teleportation can be employed — so these results may be viewed as lower bounds on achievable QEC performance. We also note that the numerical results presented here and in the appendix pertain to a single code undergoing rounds of error correction and therefore do not directly capture the fidelity of the logical $CCZ$ gate. Nevertheless, since the $CCZ$ circuit has short depth, we expect these results to serve as a reasonable proxy. A detailed simulation of the full $CCZ$ circuit is left to future work. For additional simulation details and further numerical results, see Appendix˜G.
@@
Figure 4: Circuit-level noise simulation results for tricycle codes. (a) Logical error rate for the $48,6,(d_{z}=4,d_{x}=8)]]$ code as a function of abort rate under cluster postselection (blue) and full error detection (orange). The cluster postselection data show the tradeoff between logical error rate and postselection (abort) probability using a BP+LSD decoder, while full error detection corresponds to strictly accepting only trials with no detected stabilizer flips. (b) Logical error rate versus two-qubit physical gate error rate $p_{2q}$ for $d$-rounds, fault-tolerant error correction in the $X$ basis, for tricycle codes of increasing size and distance using a most-likely error (MLE) decoder. In both panels, errors are sampled according to a standard two-qubit depolarizing circuit-level noise model and the logical error rate corresponds to the total logical error rate normalized by the number of QEC rounds and by the number of logical qubits. Logical error rates are determined via Monte Carlo simulations, with each data point corresponding to $M$ samples; error bars indicate the standard error as $\sqrt{p_{L}(1-p_{L})/M}$.